### v1

@@ 1 Introduction
In this paper, we introduce a novel compositional visual grounding method named Neuro-symbolic compositional Automaton for Visual grounding with Explicit logic Reasoning (NAVER) that addresses the limitation of existing compositional frameworks by (i) integrating explicit probabilistic logic reasoning using ProbLog [13] within a compositional pipeline to enhance visual reasoning capabilities, and (ii) incorporating a flexible, deterministic finite-state automaton (DFA) [^1] For simplicity, “automaton” refers to “deterministic finite-state automaton” in this paper. with a self-correction mechanism. Our method combines symbolic logic with large language models (LLMs) and neural-based vision foundation models, providing an explainable and logic inference for complex queries. By integrating symbolic reasoning within a DFA-based pipeline, NAVER dynamically navigates between states based on intermediate results and specific conditions, handling complex contextual constraints and relations in reasoning-based grounding. In addition, NAVER has been designed so that each step and module includes a self-correction mechanism, ensuring that mistakes occurring in one step or module, do not affect others. This combination not only improves the system’s ability to handle complex queries but also provides transparency and interpretability, thus increasing trust and facilitating error analysis. We conducted extensive experiments on popular datasets, showing that NAVER achieves state-of-the-art (SoTA) performance, comparing with SoTA end-to-end and compositional methods. These results are supported by detailed ablation studies validating the effectiveness of each component. In summary, the key contributions of this work are as follows:
@@ 2 Related Works
Recent advances in LLMs [11, 4, 35, 39] and VLMs [23, 50, 42, 62] have raised expectations for improved visual grounding in Referring Expression Detection [19] and Referring Expression Segmentation [27] tasks. The Referring Expression Detection involves identifying objects in images based on natural language descriptions, while Referring Expression Segmentation goes further by providing precise pixel-level segmentation of those objects. These approaches can be categorized into two main streams: end-to-end and (2) compositional methods.
@@
Compositional Visual Reasoning Methods. The compositional strategy offers a solution to the difficulties encountered by end-to-end methods [44, 58, 31, 17, 42]. Compositional methods address complex problems by dividing them into subtasks, solving each task independently, and then using intermediate results to solve the initial problem. By utilizing the powerful chain-of-thought (CoT) capabilities of LLMs, these methods can effectively simplify complicated challenges into smaller, solvable steps through specific instructions [17, 44]. These instructions may include execution code that includes logical operations. VisProg [17] and ViperGPT [44] reduce task-specific training by using code generation LLMs, integrating VLMs into the program to generate final output. HYDRA [20] combines a planner, RL agent, and reasoner for reasoning, outperforming other methods. It uses predefined instructions to activate perception tools and combine data for reasoning. Despite their strengths, these methods have notable limitations. They heavily rely on LLMs for reasoning, which may not fully capture complex relations like physics or geometry. Additionally, logic representations in natural language lead to ambiguity, making it hard to extract relevant information. Lastly, these systems lack self-correction mechanisms, allowing errors to propagate through the inference steps. Addressing these issues with formulated explicit logic representations and robust reasoning could enhance performance and interpretability. To address these issues, we propose NAVER, a method that improves visual reasoning by integrating symbolic logic, probabilistic logic, and a self-correcting finite-state automaton. A more detailed description of our method is provided in section 3.
@@ 3 NAVER
NAVER addresses complex visual grounding tasks that require complex reasoning, building on the limitations of previous works [17, 44, 31, 20]. In this work, we focus on referring expression detection and segmentation as the VG tasks that require reasoning with object attributes and relations among two or multiple entities within an image. Given the input $X=\{I,Q\}$ including image $I$ and a textual query $Q$, the objective of NAVER is to produce reliable and interpretable grounding result $Y$ satisfying the query $Q$. The proposed method is formulated as a finite-state automaton for a flexible pipeline, and an explicit logic reasoning ability using ProbLog [13] code generated by an LLM. To improve the robustness of the system, the full-chain of the pipeline is validated in the automaton. With this design, the system maintains a similar computational complexity to that of previous compositional methods [20], but achieves better performance (See subsection 4.3). An overview of the framework is shown in Figure 2.
@@ 4.1 Implementation Details
We implement the method using PyTorch [36] with Nvidia L40s GPUs. The vision foundation models and LLMs used in NAVER are flexible and can be exchanged for any other methods of the same type. Specifically, we use GPT-4o Mini [34] for all LLMs in the system and other compositional methods [20, 44]. We use InternVL2-8B [9] as the VLM captioner in the Perception state and as answerer in the Answering state, while XVLM [60] is used as attribute recognizer in the Logic Generation state. We also use DepthAnythingV2 [55] as depth estimator and GroundingDINO-B [30] as entity detector. For converting the bounding boxes to segmentation masks, SAM [21] is used. For segmentation tasks, we utilize GLaMM [40] for SoTA performance. For fair comparison with the previous compositional baselines [20, 44], we evaluated two versions of NAVER: (1) using the same foundation models (including GLIP [25]) as HYDRA and ViperGPT, (2) using the SoTA foundation models for best performance. We discuss the choice and the performance analysis of using different LLMs and grounding methods in subsection 4.3 and subsection 4.4. For probabilistic logic reasoning, we used Scallop [26] to enhance efficiency. The initial logic code is generated in ProbLog, then automatically translated into Scallop, a Rust-based probabilistic logic programming language optimized for high-performance inference. This translation maintains all probabilistic logic expressions while enabling faster reasoning execution than traditional ProbLog implementations, especially when dealing with large logic programs. We implement the previous baselines [20, 44, 25, 30, 10, 21, 40, 52] using the official source code, and borrow the results of other baselines from [20, 22, 61]. More details of the implementations will be elaborated in the supplementary material.
@@ 4.2 Task, Dataset and Metric
This work focuses on referring expression detection and segmentation as the two visual grounding tasks that require reasoning with object attributes and relations among multiple elements within an image. For detection, we evaluate on RefCOCO test A, RefCOCO+ test A, RefCOCOg [19] test and Ref-Adv [1] test datasets, using Intersection over Union (IoU) as the metric, following previous works [20, 44]. For the segmentation, we follow the evaluation protocol in previous works [40, 61] on the RefCOCOg and RefCLEF datasets [19] with IoU as metric. The IoU in segmentation [^2] In some previous works [22, 61, 40], it is termed as “cIoU”. is defined as the number of intersection pixels divide by the union pixels accumulated in the test set. In addition, we compare the efficiency of NAVER with other compositional baselines [20, 44] in terms of runtime, LLM token usage, and financial cost.
@@ 4.3 Quantitative Analysis
YOLO-World-X [10] 12.7 12.7 33.8 34.2
YOLO-Wrold-V2-X [10] 20.0 17.2 37.3 34.8
GroundingDINO-T [30] 48.0 48.3 52.3 52.7
GroundingDINO-B [30] 80.5 74.5 69.4 67.6
Florence2-B [52] 67.5 60.1 68.1 69.2
Florence2-L [52] 67.8 62.3 69.9 71.0
Comp. Code-bison [42] 44.4 38.2 - -
ViperGPT† [44] 59.4 58.7 60.6 58.6
HYDRA† [20] 60.4 60.1 60.8 53.5
NAVER† 67.9 63.8 61.2 63.4
ViperGPT* 64.0 68.8 63.5 59.5
HYDRA* 61.3 60.9 58.4 53.5
NAVER* 83.3 78.4 72.5 74.2
Table 2: Quantitative comparison (IoU) on Referring Expression Detection task on RefCOCO, RefCOCO+ test A set, RefCOCOg [19] and Ref-Adv [1] test set. The training set of the GroundingDINO-B and Florence2 includes RefCOCO/RefCOCO+ dataset. Other E2E methods are zero-shot. NAVER is tested with two versions, one uses the same foundation models with HYDRA and ViperGPT for fair comparison, and another one uses the SoTA models for best performance. In compositional methods, they are grouped with the same VLMs for fair comparison. The methods† and methods* use the same VFMs (GLIP-L, BLIP) and same VFMs (GroundingDINO-B, InternVL2), respectively. Both groups use GPT-4o Mini.
Referring Expression Detection. We conduct the experiments for quantitative comparison between NAVER and SoTA End-to-end (E2E) and compositional baselines on RefCOCO [19] test A, RefCOCO+ [19] test A, RefCOCOg [19] test, and Ref-Adv [1] test set. As shown in Table 2, NAVER demonstrates SoTA performance in comparison. To fairly compare to other compositional methods, NAVER use the same LLMs and foundation models (including GPT-4o Mini and GLIP [25]) as HYDRA [20] and ViperGPT [20], and performs better in all four datasets, highlighting the advantage of the design of NAVER. HYDRA, even with SoTA models, fails to achieve significant improvement due to its reliance on LLMs for generating complex Python code, which becomes a bottleneck. Using the SoTA foundation models, NAVER performs better than the SoTA E2E methods, HYDRA and ViperGPT. In the E2E methods, the training data of GroundingDINO-Base [30] and Florence2-Large [52] includes the RefCOCO and RefCOCO+, reflected in the higher IoU scores comparing to others. Notably, the YOLO-World models [10] are designed for the visual grounding without reasoning capacity, showing low IoU scores in the benchmarks.
@@
Referring Expression Segmentation. Results on RefCOCOg and RefCLEF [19] are shown in Table 3 and Table 4. In both datasets, NAVER demonstrates higher performance compared to SoTA E2E baselines and other compositional baselines [20, 44] using the same foundation models and language models. Consistent with the observation from Table 2, replacing the VFMs in HYDRA with the best performing models, i.e. GLaMM and InternVL2, incrementally improve its performance in segmentation. These experiments show that NAVER is capable of utilizing best performing VLM-based methods and yet achieve a better performance in this task.
@@
Table 5: Performance comparison with compositional methods. All results are evaluated on the RefCOCOg (IoU) test set. All methods use the same foundation models.
@@ 4.4 Ablation Studies
Contribution of Modules. To examine the contributions of key components, finite-state automaton (FSA) and logic, we conduct ablation studies on RefCOCO, RefCOCO+, RefCOCOg and Ref-Adv dataset shown in Table 7. To compare in detail the contribution of each component, we examined different combinations of the system: (1) Disabling FSA: Disabling the self-correction transitions (red arrows in Figure 2); (2) Disabling Logic: The entities detected in the Perception state are redirected to the Answerer bypassing the Logic Generation and Logic Reasoning; (3) Disabling FSA and Logic: Combining (1) and (2). Comparing rows 1 and 2, the enabling of logic leads to a significant improvement in the performance. Comparing rows 1 and 3, the FSA self-correction mechanism also provides substantial improvement. From row 4, enabling both FSA self-correction and logic improves the performance and reaches the SoTA. The impact of other modules is discussed in supplementary materials.
Selection of Grounding Models. We conduct experiments using different grounding models as the entity detector in NAVER. We tested NAVER with GLIP-Large [25], GroundingDINO-Tiny [30], GroundingDINO-Base, Florence2-Large [52], and GLaMM [40] on RefCOCO for detection and RefCOCOg for segmentation, as shown in Table 8. As NAVER is flexible in choosing foundation models, NAVER can gain performance by updating to SoTA grounding models. Comparing the performance of original grounding method with NAVER employing it, the performance is usually improved when using NAVER.
@@
FSA Logic RefCOCO RefCOCO+ RefCOCOg Ref-Adv
75.2 71.0 60.5 54.1
82.9 77.3 70.4 71.7
80.8 76.9 69.7 70.4
83.3 78.4 72.5 74.2
Table 7: Contribution of components of NAVER. Results (IoU) are based on the RefCOCO test A, RefCOCO+ test A, RefCOCOg test and Ref-Adv test set.
Method RefCOCO
GLIP-L [25] 54.1
NAVER with GLIP-L 67.9
GroundingDINO-T [30] 48.0
NAVER with GroundingDINO-T 61.0
GroundingDINO-B [30] 80.5
NAVER with GroundingDINO-B 83.3
Florence2-L [52] 67.8
NAVER with Florence2-L 70.5
@@
GLIP+SAM [25, 21] 38.4
NAVER with GLIP+SAM 54.6
GroundingDINO-B+SAM [30, 21] 61.4
NAVER with GroundingDINO-B+SAM 65.0
GLaMM [40] 74.9
NAVER with GLaMM 76.0
Table 8: Comparison of the performance using different grounding methods for NAVER. All results are evaluated on the RefCOCO (IoU) test A set and RefCOCOg (IoU) test set for referring expression detection and segmentation tasks.
@@
LLM Size Time (Sec.) IoU Time (Sec.) IoU
Deepseek-R1 [14] 70B 69.64 82.0 167.21 63.6
LLaMA3.1 [16] 70B 14.52 83.2 30.18 71.0
LLaMA3.1 [16] 8B 7.30 82.9 24.83 70.5
Gemma2 [45] 3B 6.44 83.0 19.28 69.8
GPT-4o [34] - 7.06 83.3 8.04 73.1
GPT-4o Mini [34] - 5.96 83.3 7.85 72.5
Table 9: Comparison of LLM selection for NAVER. Results are evaluated on the RefCOCO test A and RefCOCOg test set. The time values show the average seconds spent on each data sample.

### v2

@@ 1 Introduction
In this paper, we introduce a novel compositional visual grounding method named Neuro-symbolic compositional Automaton for Visual grounding with Explicit logic Reasoning (NAVER) that addresses the limitation of existing compositional frameworks by (i) integrating explicit probabilistic logic reasoning using ProbLog [14] within a compositional pipeline to enhance visual reasoning capabilities, and (ii) incorporating a flexible, deterministic finite-state automaton (DFA) [^1] For simplicity, “automaton” refers to “deterministic finite-state automaton” in this paper. with a self-correction mechanism. Our method combines symbolic logic with large language models (LLMs) and neural-based vision foundation models, providing an explainable and logic inference for complex queries. By integrating symbolic reasoning within a DFA-based pipeline, NAVER dynamically navigates between states based on intermediate results and specific conditions, handling complex contextual constraints and relations in reasoning-based grounding. In addition, NAVER has been designed so that each step and module includes a self-correction mechanism, ensuring that mistakes occurring in one step or module, do not affect others. This combination not only improves the system’s ability to handle complex queries but also provides transparency and interpretability, thus increasing trust and facilitating error analysis. We conducted extensive experiments on popular datasets, showing that NAVER achieves state-of-the-art (SoTA) performance, comparing with SoTA end-to-end [26, 31, 38, 12, 52] and compositional [43, 44, 21] methods. These results are supported by detailed ablation studies validating the effectiveness of each component. In summary, the key contributions of this work are as follows:
@@ 2 Related Works
Recent advances in LLMs [11, 4, 36, 40] and VLMs [24, 50, 43, 62] have raised expectations for improved visual grounding in Referring Expression Detection [20] and Referring Expression Segmentation [28] tasks. The Referring Expression Detection involves identifying objects in images based on natural language descriptions, while Referring Expression Segmentation goes further by providing precise pixel-level segmentation of those objects. These approaches can be categorized into two main streams: end-to-end and compositional methods.
@@
Compositional Visual Reasoning Methods. The compositional strategy offers a solution to the difficulties encountered by end-to-end methods [44, 58, 32, 18, 43]. Compositional methods address complex problems by dividing them into subtasks, solving each task independently, and then using intermediate results to solve the initial problem. By utilizing the powerful chain-of-thought (CoT) capabilities of LLMs, these methods can effectively simplify complicated challenges into smaller, solvable steps through specific instructions [18, 44]. These instructions may include execution code that includes logical operations. VisProg [18] and ViperGPT [44] reduce task-specific training by using code generation LLMs, integrating VLMs into the program to generate final output. HYDRA [21] combines a planner, RL agent, and reasoner for reasoning, outperforming other methods. It uses predefined instructions to activate perception tools and combine data for reasoning. Despite their strengths, these methods have notable limitations. They heavily rely on LLMs for reasoning, which may not fully capture complex relations like physics or geometry. Additionally, logic representations in natural language lead to ambiguity, making it hard to extract relevant information. Lastly, these systems lack self-correction mechanisms, allowing errors to propagate through the inference steps. Addressing these issues with formulated explicit logic representations and robust reasoning could enhance performance and interpretability. To address these issues, we propose NAVER, a method that improves visual reasoning by integrating symbolic logic, probabilistic logic, and a self-correcting deterministic finite-state automaton. A more detailed description of our method is provided in section 3.
@@ 3 NAVER
NAVER addresses complex visual grounding tasks that require complex reasoning, building on the limitations of previous works [18, 44, 32, 21]. In this work, we focus on referring expression detection and segmentation as the VG tasks that require reasoning with object attributes and relations among two or multiple entities within an image. Given the input $X=\{I,Q\}$ including image $I$ and a textual query $Q$, the objective of NAVER is to produce reliable and interpretable grounding result $Y$ satisfying the query $Q$. The proposed method is formulated as a deterministic finite-state automaton for a flexible pipeline, and an explicit logic reasoning ability using ProbLog [14] code generated by an LLM. To improve the robustness of the system, the full-chain of the pipeline is validated in the automaton. With this design, the system maintains a similar computational complexity to that of previous compositional methods [21], but achieves better performance (See subsection 4.3). An overview of the framework is shown in Figure 2.
@@ 4.1 Implementation Details
We implement the method using PyTorch [37] with Nvidia L40s GPUs. The vision foundation models and LLMs used in NAVER are flexible and can be exchanged for any other methods of the same type. Specifically, we use GPT-4o Mini [35] for all LLMs in the system and other compositional methods [21, 44]. We use InternVL2-8B [9] as the VLM captioner in the Perception state and as answerer in the Answering state, while XVLM [60] is used as attribute recognizer in the Logic Generation state. We also use DepthAnythingV2 [55] as depth estimator and Florence2-L [31] as entity detector. For converting the bounding boxes to segmentation masks, SAM [22] is used. For segmentation tasks, we utilize GLaMM [41] for SoTA performance. For fair comparison with the previous compositional baselines [21, 44], we evaluated two versions of NAVER: (1) using the same foundation models (including GLIP [26]) as HYDRA and ViperGPT, (2) using the SoTA foundation models for best performance. We discuss the choice and the performance analysis of using different LLMs and grounding methods in subsection 4.3 and subsection 4.4. For probabilistic logic reasoning, we used Scallop [27] to enhance efficiency. The initial logic code is generated in ProbLog, then automatically translated into Scallop, a Rust-based probabilistic logic programming language optimized for high-performance inference. This translation maintains all probabilistic logic expressions while enabling faster reasoning execution than traditional ProbLog implementations, especially when dealing with large logic programs. We implement the previous baselines [21, 44, 26, 31, 10, 22, 41, 12, 52] using the official source code, and borrow the results of other baselines from [21, 23, 61]. More details of the implementations will be elaborated in the supplementary material.
@@ 4.2 Task, Dataset and Metric
This work focuses on referring expression detection and segmentation as the two visual grounding tasks that require reasoning with object attributes and relations among multiple elements within an image. For detection, we evaluate on RefCOCO test A, RefCOCO+ test A, RefCOCOg [20] test and Ref-Adv [1] test datasets, using the most popular and widely-used metrics, i.e. accuracy [26, 31, 12, 52] and intersection over union (IoU) [44, 21] . For the segmentation, we follow the evaluation protocol in previous works [41, 61] on the RefCOCOg and RefCLEF datasets [20] with IoU as metric. The IoU in segmentation [^2] In some previous works [23, 61, 41], it is termed as “cIoU”. is defined as the number of intersection pixels divide by the union pixels accumulated in the test set. In addition, we compare the efficiency of NAVER with other compositional baselines [21, 44] in terms of runtime, LLM token usage, and financial cost.
@@ 4.3 Quantitative Analysis
YOLO-World-X [10] 12.1 12.1 32.9 32.2
YOLO-Wrold-V2-X [10] 19.8 16.8 36.5 33.1
GroundingDINO-T [31] 61.6 59.7 60.6 60.5
GroundingDINO-B [31] 90.8 84.6 80.3 78.0
SimVG [12] 94.9 91.0 88.9 74.4
Florence2-B [52] 94.5 91.2 88.3 72.2
Florence2-L [52] 95.1 92.5 90.9 71.8
Compositional Code-bison [43] 44.4 38.2 - -
ViperGPT† [44] 62.6 62.3 67.2 60.7
HYDRA† [21] 60.9 56.5 62.9 54.4
NAVER† 70.1 64.1 69.5 65.1
ViperGPT* 68.6 73.8 68.7 58.2
HYDRA* 65.7 66.2 59.9 48.3
NAVER* 96.2 92.8 91.6 75.4
Table 2: Quantitative comparison (accuracy) on referring expression detection task on RefCOCO, RefCOCO+, RefCOCOg [20] and Ref-Adv [1] set. The training set of the GroundingDINO-B, SimVG and Florence2 includes RefCOCO/RefCOCO+/RefCOCOg datasets. Other E2E methods are zero-shot. NAVER is tested with two versions, one uses the same foundation models with HYDRA and ViperGPT for fair comparison, and another one uses the SoTA models for best performance. In compositional methods, they are grouped with the same VLMs for fair comparison. The methods† and methods* use the same VFMs (GLIP-L, BLIP) and same VFMs (Florence2-L, InternVL2), respectively. Both groups use GPT-4o Mini.
Referring Expression Detection. We conduct the experiments for quantitative comparison between NAVER and SoTA end-to-end (E2E) and compositional baselines on RefCOCO [20] test A, RefCOCO+ [20] test A, RefCOCOg [20] test, and Ref-Adv [1] test set. As shown in Table 2, NAVER demonstrates SoTA performance in comparison. To fairly compare to other compositional methods, NAVER use the same LLMs and foundation models (including GPT-4o Mini and GLIP [26]) as HYDRA [21] and ViperGPT [21], and performs better in all four datasets, highlighting the advantage of the design of NAVER. ViperGPT and HYDRA, even with SoTA models, fail to achieve significant improvement due to its reliance on LLMs for generating complex Python code, which becomes a bottleneck. Using the SoTA foundation models, NAVER performs better than the SoTA E2E methods, as well as compositional approaches such as HYDRA and ViperGPT, because NAVER is equipped with self-correction mechanism and logic reasoning, also breaks down the complex reasoning into simpler subtasks, which are easier to be solved by LLMs and VLMs. For more comprehensive results and comparison, please see Table 10 in supplementary material including more methods and metrics.
@@
Referring Expression Segmentation. Results on RefCOCOg and RefCLEF [20] are shown in Table 3 and Table 4. In both datasets, NAVER demonstrates higher performance compared to SoTA E2E baselines and other compositional baselines [21, 44] using the same foundation models and language models. Consistent with the observation from Table 2, replacing the VFMs in compositional methods with the best performing models, i.e. GLaMM and InternVL2, incrementally improve their performances in segmentation. These experiments show that NAVER is capable of utilizing best performing VLM-based methods and yet achieve a better performance in this task.
@@
Table 5: Performance comparison with compositional methods. All results are evaluated on the RefCOCOg (IoU) test set for segmentation. All methods use the same foundation models.
@@ 4.4 Ablation Studies
Contribution of Modules. To examine the contributions of key components, deterministic finite-state automaton (DFA), logic and answerer, we conduct ablation studies on RefCOCO, RefCOCO+, RefCOCOg and Ref-Adv dataset shown in Table 7. To compare in detail the contribution of each component, we examined different combinations of the system: (1) Disabling DFA: Disabling the self-correction transitions (red arrows in Figure 2); (2) Disabling Logic: The entities detected in the Perception state are redirected to the Answerer bypassing the Logic Generation and Logic Reasoning; (3) Disabling Answerer: the answer is not validated by the answerer; and the feasible combinations of (1), (2) and (3). Comparing rows 1 and 4, 3 and 6, 5 and 7, the DFA self-correction mechanism provides substantial performance improvement. Comparing rows 1 and 3, 2 and 5, the enabling of logic leads to a significant improvement in the performance. Comparing rows 1 and 2, 3 and 5, 6 and 7, the answerer validates the result and improves the accuracy. From row 7, enabling all DFA self-correction, logic and answerer improves the performance and reaches the SoTA.
Selection of Grounding Models. We conduct experiments using different grounding models as the entity detector in NAVER. We tested NAVER with various grounding methods [26, 31, 12, 41] on RefCOCO for detection and RefCOCOg for segmentation, as shown in Table 8. As NAVER is flexible in choosing foundation models, NAVER can gain performance by updating to SoTA grounding models. Comparing the performance of original grounding method with NAVER employing it, the performance is usually improved when using NAVER.
DFA Logic Answerer Ref Ref+ Refg Ref-Adv
51.7 51.3 48.8 34.1
57.4 57.6 50.8 34.3
57.8 62.0 51.3 36.4
68.5 69.3 70.1 64.7
78.1 68.4 53.4 43.6
76.2 73.8 75.8 69.7
96.2 92.8 91.6 75.4
Table 7: Contribution of components of NAVER. Results (accuracy) are based on the RefCOCO test A, RefCOCO+ test A, RefCOCOg test and Ref-Adv test set.
@@
RefCOCO
Method Itself In NAVER
GLIP-L [26] 55.0 70.1
GroundingDINO-T [31] 61.6 64.9
GroundingDINO-B [31] 90.8 91.7
SimVG [12] 94.9 95.4
Florence2-L [52] 95.1 96.2
@@
GLIP+SAM [26, 22] 38.4 54.6
GroundingDINO-B+SAM [31, 22] 61.4 65.0
GLaMM [41] 74.9 76.0
Table 8: Comparison of the performance using different grounding methods for NAVER. All results are evaluated on the RefCOCO (accuracy) for detection and RefCOCOg (IoU) for segmentation. The column Itself shows the performance of the methods, and column In NAVER shows the performance of NAVER using these methods for grounding.
@@
LLM Size Time (Sec.) Acc. Time (Sec.) Acc.
DeepSeek-R1 [15] 70B 71.27 95.4 82.13 85.8
LLaMA3.1 [17] 70B 12.41 96.0 19.40 88.8
LLaMA3.1 [17] 8B 8.23 95.0 9.25 90.5
Gemma2 [45] 3B 8.18 96.1 9.13 91.1
GPT-4o [35] - 8.67 96.6 10.29 89.7
GPT-4o Mini [35] - 5.74 96.2 7.91 91.6
Table 9: Comparison of LLM selection for NAVER. Results are evaluated on the RefCOCO and RefCOCOg for detection. The time values show the average seconds spent on each data sample.